# Near-Optimal Approximation of Matrix Functions by Lanczos-FAN. Amsel, T. Chen, A. Greenbaum, C. Musco, and C. Musco

### v2

@@ A.1 Instance, spectrum, and FOV optimality can coincide
First, with high probability, $\|\mathbf{b}\|_{2}=\Theta(\sqrt{d})$; see, for example, [Ver18, 3.1.2]. Second, we show that with high probability $|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{b}|=\Omega(1/d)$ for all $j$. Since $\mathbf{V}$ is orthonormal, $\mathbf{w}:=\mathbf{V}^{\mathsf{T}}\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. By anti-concentration of the normal distribution, the probability that $|w_{j}|>\epsilon$ is at least $1-0.4\epsilon$ for any $\epsilon$. By a union bound, this holds simultaneously for all $j$ with probability at least $1-0.4\epsilon d$. Setting $\epsilon=\Theta(1/d)$ finishes the argument. Finally, by another union bound, our bounds on the numerator and on the denominator hold simultaneously with high probability.
A.3 Lanczos-FA on indefinite systems
The MINRES iterates are defined as
$\hat{\mathbf{y}}_{k}:=\operatornamewithlimits{argmin}_{\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{b}-\mathbf{A}\mathbf{y}\|_{2}=\operatornamewithlimits{argmin}_{\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{y}\|_{\mathbf{A}^{2}}.$
Define the residual vectors
$\mathbf{r}_{k}:=\mathbf{b}-\mathbf{A}\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b}),\qquad\mathbf{r}_{k}^{\mathrm{M}}:=\mathbf{b}-\mathbf{A}\hat{\mathbf{y}}_{k},$
and note that the MINRES residual norms are non-increasing due to the optimality of the MINRES iterates. In [CG96], it is shown that the CG residual norms are near the MINRES residual norms at iterations where MINRES makes good progress. More precisely, the algorithms are related by
$\|\mathbf{r}_{k}\|_{2}=\frac{\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\sqrt{1-\left(\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}/\|\mathbf{r}_{k-1}^{\mathrm{M}}\|_{2}\right)^{2}}}.$
In [CGMM22, Appendix A], this result was used to show that Lanczos-FA has an effective geometric convergence rate. The following theorem, which we stated as Eq. 7, represents a strengthening of [CGMM22, Appendix A]. To the best of our knowledge, the following result is new.
Theorem A.4.
For every $k>0$, there exists $k^{*}\leq k$ such that
${\|\mathbf{b}-\mathbf{A}\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}}\leq\left(\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}\right)\min_{\deg(p)<k}{\|\mathbf{b}-\mathbf{A}p(\mathbf{A})\mathbf{b}\|_{2}}.$
Switching from the 2-norm of the residual ($\mathbf{A}^{2}$-norm of the error) to the 2-norm of the error gives the bound stated in Eq. 7.
Proof A.5.
Fix $k>0$. If $\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}>\|\mathbf{r}_{0}^{\mathrm{M}}\|_{2}=\|\mathbf{b}\|_{2}$, then the theorem holds with $k^{*}=0$ as $\sqrt{\mathrm{e}}\leq(\mathrm{e}\sqrt{k}+1/\sqrt{k})$. Assume $\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}\leq\|\mathbf{r}_{0}^{\mathrm{M}}\|_{2}=\|\mathbf{b}\|_{2}$. Then there exists an integer $k^{\prime}\in[0,k)$ such that
$\|\mathbf{r}_{k^{\prime}}^{\mathrm{M}}\|_{2}\geq\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}\geq\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}.$
This implies that
$\frac{1}{\sqrt{\mathrm{e}}}\geq\frac{\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{\prime}}^{\mathrm{M}}\|_{2}}=\frac{\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{\prime}}^{\mathrm{M}}\|_{2}}\cdot\frac{\|\mathbf{r}_{k^{\prime}+2}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}}\>\cdots\>\frac{\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k-1}^{\mathrm{M}}\|_{2}}.$
Thus, there exists an integer $k^{*}\in(k^{\prime},k]$ such that
$\frac{\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{*}-1}^{\mathrm{M}}\|_{2}}\leq\left(\frac{1}{\sqrt{\mathrm{e}}}\right)^{1/(k-k^{\prime})}\leq\left(\frac{1}{\sqrt{\mathrm{e}}}\right)^{1/k}.$
Since the MINRES residual norms are non-increasing, $\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}\leq\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}\leq\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}$. Thus,
$\|\mathbf{r}_{k^{*}}\|_{2}=\frac{\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}}{\sqrt{1-\left(\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}/\|\mathbf{r}_{k^{*}-1}^{\mathrm{M}}\|_{2}\right)^{2}}}\leq\frac{\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\sqrt{1-(1/\sqrt{\mathrm{e}})^{2/k}}}.$
It can be verified that
$\mathrm{e}\sqrt{k}\leq\frac{\mathrm{e}}{\sqrt{1-(1/\mathrm{e})^{1/k}}}\leq\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}.$
Thus, we find that
$\|\mathbf{b}-\mathbf{A}\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}=\|\mathbf{r}_{k^{*}}\|_{2}$ $\leq\left(\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}\right)\|\mathbf{r}_{k}^{\textup{M}}\|_{2}.$
The result follows from the optimality of the MINRES iterates.

### v3

@@ C.1 Relation between spectrum and instance optimality for random vectors
First, with high probability, $\|\mathbf{b}\|_{2}=\Theta(\sqrt{d})$; see, for example, [vershynin_2018, 3.1.2]. Second, we show that with high probability $|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{b}|=\Omega(1/d)$ for all $j$. Since $\mathbf{V}$ is orthonormal, $\mathbf{w}:=\mathbf{V}^{\mathsf{T}}\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. By anti-concentration of the normal distribution, the probability that $|w_{j}|>\epsilon$ is at least $1-0.4\epsilon$ for any $\epsilon$. By a union bound, this holds simultaneously for all $j$ with probability at least $1-0.4\epsilon d$. Setting $\epsilon=\Theta(1/d)$ finishes the argument. Finally, by another union bound, our bounds on the numerator and on the denominator hold simultaneously with high probability.
Appendix D Proofs of Theorem 6 and Theorem 7
We begin by quoting bounds from [chen_greenbaum_musco_musco_22]:
Lemma 16.
For all $k\geq 1$, the Lanczos-FA iterate satisfies the bounds
  1. 1.
  2. 2.
Proof.
For Item 1, see Example 4.1 in [chen_greenbaum_musco_musco_22]. For part Item 2, we use the same proof again but substituting the inverse square root for the square root. For both parts, we have simplified the resulting bounds using the fact that for $k\geq 1$,
Lemma 17.
For all $k\geq 1$, the Lanczos-FA iterate satisfies the bounds
Proof.
Let
and note that $p(x)^{2}$ is a polynomial of degree less than $k$. Therefore,
Thus, since $|1/x-p^{*}(x)^{2}|=|1/\sqrt{x}+p^{*}(x)|\cdot|1/\sqrt{x}-p^{*}(x)|$, we can plug the above into Equation 17 to obtain a bound
Substituting this inequality above proves the lemma. ∎
The main proofs now follow directly.
Proof of Theorem 6.
Substitute Lemma 17 into Lemma 16, part (b). ∎
Proof of Theorem 7.
Let
and note that $\tilde{p}$ is a polynomial of degree less than $k/2$. Then
Combining Lemma 16 part (a), Lemma 17, and the above proves the theorem. ∎
Appendix E Additional Experiments
E.1 Validating Theorems 6 and 7
To understand Theorems 6 and 7, we repeat the experiment using the inverse square root and square root functions. We reuse the second and third $\mathbf{A}$ matrix and $\mathbf{b}$ vectors from the previous experiment. As Figure 7 shows, our bounds are tighter than Fact 3. They closely resemble the true convergence of Lanczos-FA, but stretched horizontally by a factor of 2. This is because the degree of the polynomial minimization in both bounds is $k/2$, while in these examples Lanczos-FA performs nearly instance optimally (that is, as well as the best degree $k$ polynomial). Still, our bounds capture the correct qualitative behavior of the algorithm.
Figure 7: The bounds of Theorems 6 and 7 capture the convergence behavior of Lanczos-FA for $\mathbf{A}^{\pm 1/2}\mathbf{b}$. In particular, they can be tighter than the FOV optimality bound of Fact 3. The predicted rate of convergence is about half that observed for Lanczos-FA in this example; this is due to the parameter $c=1/2$ in both bounds.
E.2 Rational functions with poles between eigenvalues
In Theorem 4, we assume that the poles of the rational function $r(x)$ lie outside the interval $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ containing $\mathbf{A}$’s eigenvalues. If the poles of $r(x)$ live in $\mathcal{I}$, then the Lanczos-FA iterate can be arbitrarily far from the optimal iterate. Indeed, an eigenvalue of $\mathbf{T}$ might be very close to a pole, causing $f(\mathbf{T})$ to be poorly behaved. This behavior is easily observed e.g. for $f(\mathbf{A})=\mathbf{A}^{-1}$ and rules out a direct analog of Theorem 4 when $r(x)$ has poles in $\mathcal{I}$. However, as noted in Equation 18, for $1/x$ there exists bounds for Lanczos-FA in terms of the best possible KSM [chen_meurant_24].
However, as discussed in Section B.1, the “overall” convergence of Lanczos-FA still seems to closely follow the optimal approximation for many matrix functions. Specifically, even if the Lanczos-FA iterate deviates significantly from optimal on some iterations, there are other iterations where it is very close to optimal. We illustrate this phenomena in Figure 6 for two rational functions with poles in $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, as well as the function $\operatorname{sign}(x)$, which has a discontinuity in $\mathcal{I}$ and would typically be approximated by a rational function with conjugate pairs of imaginary poles on the imaginary axis which become increasingly close to the the interval as the rational function degree increases. [^5] We remark that [chen_greenbaum_musco_musco_23] shows that applying their Lanczos-OR method to each term in the partial fraction decomposition of the approximating rational function yields an approximation to the sign function that seems to avoid the oscillations of Lanczos-FA.
E.3 Comparison with Lanczos-OR
As discussed in Section B.2, the Lanczos-OR method is an alternative method for rational matrix function approximation for which a tighter guarantee than Theorem 4 is currently known (the iterates are optimal in a certain norm). However, Lanczos-FA usually outperforms Lanczos-OR when measured in the 2-norm. For example, as discussed in Section 4.1, when approximating $\mathbf{A}^{-q}\mathbf{b}$, the worst optimality ratio of Lanczos-FA that we were able to observe was $O(\sqrt{q\cdot\kappa(\mathbf{A})})$, much lower than Theorem 4 would suggest). Repeating the same experiment with Lanczos-OR shows that the method’s optimality ratio appears to grow as $\Omega(\kappa(\mathbf{A})^{q/2})$ (see Figure 8).
Figure 8: The maximum observed ratio between the error of Lanczos-OR and the optimal error over choice of right hand side $\mathbf{b}$ when approximating $\mathbf{A}^{-q}$ for matrices with varying condition number $\kappa$. Each point shows the optimality ratio for a different pair of $\kappa$ and $q$. Points with the same color correspond to the same value of $\kappa$. On the left, the dotted line plots $g(q)=\kappa^{q/2}$ for the maximum $\kappa$ considered ($10^{6}$). On the right, the dotted line plots $g(\kappa)=\kappa^{q/2}$ for the maximum $q$ considered ($642^{6}=$ 64). Overall, the optimality ratio appears to grow as $\Omega(\kappa^{q/2})$. Contrast with Figure 3.